# How Much Do LLMs Hallucinate across Languages? On Multilingual Estimation of LLM Hallucination in the Wild

### v3

@@ 5 Conclusion
We presented the first effort towards understanding how much multilingual LLMs hallucinate “in the wild”. To this end, we proposed a novel framework for hallucination rate estimation, which adjusts the number of detected hallucinations based on the detector’s performance resulting in more reliable rate estimates. We trained a series of multilingual detection models, and measured their precision and recall scores on our newly created mFAVA datasets across 30 languages. To estimate hallucinations, we build a novel synthetic open-domain knowledge-intensive QA dataset for which we collected answers from eleven open-source LLMs. Our findings indicate that smaller models and models that cover more languages hallucinate significantly more, and that model response-length does not correlate with hallucination rate.
@@
Finally, we deliberately limited the scope of this work to assessing factual correctness and we do not cover factual coverage. We decided to do so as quantifying hallucinations in long-form generation is already difficult for English Xu et al. (2023); Min et al. (2023); Wei et al. (2024b) and more so in non-english languages (Kim et al., 2024), and
@@ Acknowledgments
This work was supported by the Alcatel-Lucent Stiftung and Deutsches Stiftungszentrum through the grant “Equitably Fair and Trustworthy Language Technology” (EQUIFAIR, Grant Nr. T0067/43110/23). The authors gratefully acknowledge the computing time granted by the John von Neumann Institute for Computing (NIC) and provided on the supercomputer JURECA (Jülich Supercomputing Centre, 2021) at Jülich Supercomputing Centre (JSC).
@@ A.5 Hallucination Evaluation Dataset
We use Wikipedia article summary (text before the first heading) as references and prompt gpt-4 to generate 2 knowledge-intensive queries per article. Sometimes, it generated only one query even though we explicitly state to generate two queries. We did not prompt the GPT-4 again to generate the second query due to budget constraints. The total cost to generate prompts for 31 languages is \$192. Per language statistics can be found in Table 9. We will release hallucination evaluation data under an open scientific licensing.

### v4

@@ 5 Conclusion
We presented the first effort towards understanding how much multilingual LLMs hallucinate in a realistic setting. To this end, we proposed a novel framework for hallucination rate estimation, which adjusts the number of detected hallucinations based on the detector’s performance resulting in more reliable rate estimates. We trained a series of multilingual detection models, and measured their precision and recall scores on our newly created mFAVA datasets across 30 languages. To estimate hallucinations, we build a novel synthetic open-domain knowledge-intensive QA dataset for which we collected answers from eleven open-source LLMs. Our findings indicate that smaller models and models that cover more languages hallucinate significantly more, and that model response-length does not correlate with hallucination rate. Lastly, linear mixed effects analysis reveals that these factors do not operate independently—model size acts as a critical moderator, with smaller models suffering disproportionately when supporting multiple languages and generating longer responses.
@@
Finally, we deliberately limited the scope of this work to assessing factual correctness and we do not cover factual coverage. We decided to do so as quantifying hallucinations in long-form generation is already difficult for English Xu et al. (2023); Min et al. (2023); Wei et al. (2024b) and more so in non-english languages (Kim et al., 2024), and currently, resources for assessing multilingual factual coverage are still lacking.
@@ Acknowledgments
This work was supported by the Alcatel-Lucent Stiftung and Deutsches Stiftungszentrum through the grant “Equitably Fair and Trustworthy Language Technology” (EQUIFAIR, Grant Nr. T0067/43110/23). The work of Anne Lauscher is funded under the Excellence Strategy of the German Federal Government and States. The authors gratefully acknowledge the computing time granted by the John von Neumann Institute for Computing (NIC) and provided on the supercomputer JURECA (Jülich Supercomputing Centre, 2021) at Jülich Supercomputing Centre (JSC).
@@ A.5 Hallucination Evaluation Dataset
We use Wikipedia article summary (text before the first heading) as references and prompt gpt-4 to generate 2 knowledge-intensive queries per article. Sometimes, it generated only one query even though we explicitly state to generate two queries. We did not prompt the GPT-4 again to generate the second query due to budget constraints. The total cost to generate prompts for 31 languages is \$192. Per language statistics can be found in Table 9. We release hallucination evaluation data under an open scientific licensing.